OpenStudy (anonymous):

Determine the area bounded by the curve=cos x, the x-axis, and the vertical x=π and x=3π/2.

OpenStudy (anonymous):

Do you have any ideas how to start?

OpenStudy (anonymous):

Remember that the Definite integral from a to b of a function, represents the area under that function.

OpenStudy (anonymous):

\[\int\limits_a^b f(x) dx = Area\]

OpenStudy (anonymous):

I think it's \[\int\limits_{\pi /4}^{3\pi/2}\cos x dx\]. Is it correct?

OpenStudy (anonymous):

It's almost correct. The lower limit is not pi/4

OpenStudy (anonymous):

It is just pi

OpenStudy (anonymous):

Oh yeah! Thank you very much!

OpenStudy (anonymous):

U r welcome